OpenStudy (anonymous):

what is the slope of a line that is perpendicular to the line represented by the equation 3x+4y=12? a)4/3 b)3/4 c)-3/4 d)-4/3

OpenStudy (aroub):

a

OpenStudy (anonymous):

give x = 0 and ten y = 0 : 0 + 4y= 12 ..... y = 3 3x + 0 = 12 .... x = 4 then the plots coordinate will be x= 4 y = 3 soo the slope = tanx = y/x = 3 / 4

OpenStudy (anonymous):

aroub you solved wrong! be carefull

OpenStudy (aroub):

why?

OpenStudy (aroub):

its correct

OpenStudy (anonymous):

look slope is height / width... you did width / height here y is the height and x is the width and x = 4 ,y = 3 in here so it will be height / width = y / x = 3 / 4 ok?

OpenStudy (aroub):

no, y=mx+b 4y=-3x+12 y=-3/4x+3 the slope perpendicular =4/3

OpenStudy (anonymous):

well to use this equation you need the coordinates of the plot... so firlst give y= 0 x x= 0 then find the plot, then draw a line (you willsee the lines slope willbe equal to this line) so you will see that its y/x !!

OpenStudy (aroub):

korcan i dont think so

OpenStudy (anonymous):

aroub i don'T think so

OpenStudy (anonymous):

ask to amistre...

OpenStudy (aroub):

hahaha ok :)

OpenStudy (anonymous):

The slope of the line perpendicular to the given equation is 4/3. Since the slope of the line is -3/4, to find the perpendicular slope, multiply by -1 and reverse the denominator and numerator.

OpenStudy (anonymous):

required slope = -1/given slope = -1 over -3/4 = 4/3.